OpenStudy (anonymous):

absolute value 16=5|p|-4

TheSmartOne (thesmartone):

Well first thing we can do it add 4 on both sides :)

OpenStudy (anonymous):

is this the complete equation that you have typed out?

TheSmartOne (thesmartone):

And then divide by 5 on both sides.

TheSmartOne (thesmartone):

And \(\sf |p|=x\) Then \(\sf -p = x\\ ~~p=x\)

OpenStudy (jdoe0001):

hmmm actually should be +4... anyhow lemme fix that quick

OpenStudy (jdoe0001):

hmmm

OpenStudy (jdoe0001):

hh shoot.. lemme fix that too

OpenStudy (jdoe0001):

\(\large { 16=5|p|-4\implies 16+4=5|p|\implies \cfrac{16+4}{5}=|p| \\ \quad \\ \begin{array}{llll} \cfrac{16+4}{5}=-(p)\implies ?\\ \\\hline\\ \cfrac{16+4}{5}=+(p)\implies ? \end{array} }\)